OpenStudy (milo123):

The height of a cone-shaped container is 18 centimeters and its radius is 7 centimeters. Meg fills the container completely with honey. Every day Meg uses 21 cubic centimeters of honey from the container to make honey cookies. After how many days will the container be completely empty? Round your answer to the nearest whole number. (Use π = 3.14.) 44 66 78 92

OpenStudy (rcp031):

First Find the volume of a cone-shaped container=1/3*(3.14*r^2*h) here r=radius h=height Then divide total volume by 21cc to days will the container be completely empty

563blackghost (563blackghost):

Lets first find the volume of the cone :) So lets write out the formula for it... \(\huge\bf{V=\pi r^{2} \frac{h}{3}}\) We know that the height is `18cm` and the radius is `7` lets input that.... \(\huge\bf{V = 3.14(7^{2}) \frac{18}{3}}\) First simplify the parenthesis... \(\huge\bf{V = 3.14(49) \frac{18}{3}}\) Multiply... \(\huge\bf{V=(153.86)\frac{18}{3}}\) Then divide the fraction... \(\huge\bf{V=(153.86) 6}\) Then multiply to find the volume..What would the volume equal?

OpenStudy (milo123):

923.16 ? @563blackghost

OpenStudy (milo123):

is it 66?

563blackghost (563blackghost):

Not quite. Sorry I was offline :(